# Advancing Time Series Wildfire Spread Prediction: Modeling Improvements and the WSTS+ Benchmark

### v2

@@ 6.2 Time-Series Input, $T=5$
Existing models for the time-series scenario generally adopt one of two approaches: (i) a data-level fusion, or (ii) a feature-level fusion. In data-level fusion, the features for each day, $\tilde{x}(t)\in\mathbb{R}^{H\times W\times C}$ are concatenated along the feature dimension into one input tensor, $x(t)=[\tilde{x}(t-1)|,...,|\tilde{x}(t-T)]\in\mathbb{R}^{H\times W\times CT}$, after which they can be processed in the same manner as single-day input (see Sec. 2 for problem notation). Therefore, we adopt our best-performing $T=1$ models from Sec. 6.1, and their hyperparameter settings, and evaluate them for data-level fusion. As a reference, we also include the reported results of the Res18-Unet [15] when it was applied for data-level fusion.
In this context, feature-level fusion implies that we use a shared encoder to first extract features (or embeddings) independently for each day of our input, $\tilde{z}(t)=f_{\theta_{En}}(\tilde{x}(t))$ so that we have a collection of features, $z(t)=\{\tilde{z}(t-i)\}_{i=1}^{T}$, which are utilized as input into a subsequent model for joint processing (i.e., fusion). The current SOTA accuracy on WSTS, both for the time-series setting, and overall, was obtained with a UTAE model [14], as reported in [15]. Furthermore, the UTAE achieved superior accuracy despite having just 1.1M parameters - significantly fewer than many other models considered (e.g., the Res18-Unet has 14.3M). Therefore, we focus our modeling improvements on the UTAE from [15].
@@
We develop two improved UTAE models, referred to as UTAE and UTAE(Res18). We discuss the design of each model next.
Our UTAE Model. Our UTAE includes two major improvements over the UTAE[15]. The first improvement is to adopt all of the changes investigated for the single-day models from Sec. 6.1. Pursuant to this, following previous work convention, we did a joint search over the following hyperparameters using a single fold of the WSTS benchmark: pre-training (or not), learning rates ($[1e-2,1e-3,1e-4,1e-5]$), and the type of loss (Focal, BCE, Jaccard, and Dice loss). The second improvement is the introduction of relative positional encodings in the temporal fusion utilized by the UTAE, to replace the day-of-year embeddings (see Sec. 5 for review of topic, and notation). Specifically, instead of using positional encodings as in [14, 15] that are based upon the absolute day of the year, where $\bar{t}\in[1,365]$, we propose to use a variant of relative positional encoding that indicates where in the daily sequence each feature is located, so that $\bar{t}\in[1,...,T]$ for a T-day input. We hypothesize that the features (especially the fire mask) from the most recent day of the fire will be most important for making predictions, and therefore the relative position of each feature embedding in the sequence is much more important than its position in the year. Furthermore, it may be difficult for the models to infer relative positional information from absolute encodings, potentially undermining performance.
Our UTAE(Res18) Model. This model is obtained by making one additional improvement to our UTAE model. The encoder utilized in the UTAE[15] is relatively small (in terms of free parameters). Therefore, in similar fashion to our investigation in Sec. 6.1, we replace the existing UTAE’s encoder with a pre-trained ResNet-18.
@@
Tab. 2 reports the accuracy (in terms of AP) of our time-series models on the WSTS benchmark, categorized by the type of fusion performed: data-level or feature-level. Regarding data-level fusion, our Res18-Unet and Swin-Unet both substantially outperform the existing Res18-Unet[15] across all combinations of input features, with the Res18-Unet providing the best overall AP (AP=0.472, on Vegetation features). Regarding feature-level fusion, our two UTAE models (UTAE and UTAE(Res18)) substantially outperform the existing UTAE[15], which is the current SOTA model on WSTS, both for time-series input ($T>1$), and overall. Our UTAE(Res18) model achieves the highest overall performance for each combination of input features, across both single-day and time-series models. In particular, our UTAE(Res18) achieves the highest overall AP with the Vegetation (Veg) feature subset, leading to a new overall SOTA performance on WSTS of AP=0.478.
Notably, our results indicate that models receiving time-series input generally outperform those with single-day input. This is especially apparent when comparing data-level fusion models, such as Res18-Unet and SwinUnet, with their single-day counterparts, since they have few architectural differences. Most existing wildfire spread prediction in the literature has focused on the single-day input, however our findings here corroborate those from [15], and suggest that time-series modeling is a promising emerging modeling strategy.
Our results also provide evidence that each of our modeling improvements is beneficial. As discussed, our UTAE included several applicable improvements discussed for our single-day models in Sec. 6.1, as well as our improved temporal encodings described in this sub-section. We therefore conducted an ablation experiment, reported in Tab. 3, to demonstrate that our modified positional encodings provide additional benefits. To demonstrate that the pre-trained ResNet-18 encoder is beneficial, we can compare the performance of UTAE(Res18) and UTAE in Tab. 2: the pre-trained ResNet-18 is the only difference between these two models.
Table 3: Test AP of UTAE trained on Vegetation features using the original Absolute positional encodings from [15], versus our proposed Relative positional encodings
@@ 7 The WSTS+ Benchmark
Our results on the WSTS benchmark indicated that relatively simple models performed best, such as those based upon a ResNet-18, rather than models utilizing larger encoders (e.g., ResNet-50) or those utilizing attention (e.g., SwinUnet). This contrasts sharply with the broader vision literature where larger models tend to perform best, given sufficient quantities of training data. Therefore, we hypothesize that collecting more training data would facilitate the use of larger models, yielding superior modeling performance. To investigate this hypothesis, we expand the original WSTS benchmark by curating four additional years of historical wildfire data: 2016, 2017, 2022, and 2023. Our extended dataset, termed WSTS+, contains twice the number of years of historical wildfire data, expands the geographic diversity of the benchmark, and is – to our knowledge – the largest public benchmark for time-series next-day wildifre spread prediction. We visualize the geographic distribution of WSTS+ events in Fig. 4 and observe that it is mostly similar to WSTS around the Western United States, with some distinction in more eastern states. Tab. 4 summarizes the differences between both datasets in terms of numbers of years, fire events, total images, and active fire pixels. Further collection details can be found in the Supplement.
@@
Years 4 (2018-2021) 8 (2016-2023) $+100$
Fire Events 607 1,005 $+65.6$
Total Images 13,607 24,462 $+79.8$
Active Fire Px 1,878,679 2,638,537 $+40.4$
@@ 7.1 Benchmarking Models with WSTS+
As compared to WSTS, we propose a new scheme for evaluating models using WSTS+, which exploits its greater size to significantly reduce computational complexity compared to WSTS – thereby making the benchmark more accessible to researchers – while maintaining a similar level of real-world rigor. For WSTS+, we propose to divide the available data into four folds that each contain two consecutive years of historical wildfire data. We then evaluate models using four-fold cross-validation, where in each iteration, one fold of data is used for testing, one fold for validation, and two folds for training, as illustrated in Fig. 5. To ensure that the testing and validation sets have the same relative temporal distance to the training set, we always select them so that they are non-consecutive. This results in four-fold cross-validation instead of the twelve-fold cross-validation utilized in WSTS, making it much more accessible to researchers. At the same time, this approach doubles the quantity of data in the training and validation sets, ideally allowing researchers to train larger and more accurate models. Lastly, because two consecutive years of data are included in the test set, the benchmark still evaluates models under highly realistic and rigorous testing conditions,
@@
Res18-Unet-LTAE 5 $\mathbf{0.354\pm 0.113}$ $\mathbf{0.363\pm 0.129}$ ${0.350\pm 0.117}$
@@ 7.2 Experimental Results with WSTS+
Using our updated cross-validation scheme, we train our best $T=1$ models and our best $T=5$ model on WSTS+ and report the results in terms of mean average precision across all three feature sets in Tab. 5. We see that the performance rank-order of our three models is still similar on WSTS+ as compared to WSTS. However, the overall performance is significantly lower for these models on WSTS+ as compared to WSTS (by roughly 0.1 AP), and the differences in AP between the models are lower. These results seem to contradict our initial hypothesis that more training data would enable larger models and improve accuracy.
To investigate further, in Fig. 6, we measure the performance (in AP) across all eight years of data of our Res18-Unet model trained on WSTS or WSTS+, respectively. To estimate the WSTS+ models, we simply report the cross-validation test performance, stratified by year. We apply the same approach to the WSTS benchmark, obtaining performance metrics for 2018 to 2021. We select the best-performing Res18-Unet (as judged by its test fold error) and evaluate it to the four new years of WSTS data. The results in Fig. 6 indicate that, when viewed this way, the performances of the WSTS and WSTS+ models are very similar. In particular, models trained on each benchmark obtain similar AP on the original four years of the WSTS benchmark (the bolded years in Fig. 6), indicating that the two additional years of data that were available in each training fold of WSTS+ were not harmful; however, they were not helpful either, contradicting our initial hypothesis. Fig. 6 also reveals that the overall difficulty of the new years of data - 2016, 2017, and 2022 in particular - are higher, as evidenced by the fact that all models perform much worse on those years, explaining why the overall AP on the WSTS+ is lower than WSTS.
Crucially, the results in Fig. 6 reveal substantial variance in model prediction accuracy across the eight years in WSTS+, regardless of which of our two models were utilized: WSTS or WSTS+. A similar pattern was observed on the WSTS benchmark [15], where the authors found similar variation in performance, which was stable across models trained with different years of training data. These findings could admit several explanations: label noise variations across different years, making some harder than others; each year may exhibit different feature distributions (i.e., domains shifts), or wildfire behavior may vary across years (i.e., concept shifts). Fig. 7 presents a UMAP visualization of the input features of WSTS years (in blue) and the new WSTS+ years (in red) at the deepest level of our best Res18-Unet encoder. While there is substantial overlap in the features obtained from each year of data, we also observe that many years seem to exhibit unique feature patterns. If some years exhibit especially unique feature distributions that are not present in any other years (or few other years) then it could contribute to low accuracy on those years.
An Emerging Wildfire Modeling Challenge?
Our results suggest that naively combining historical wildfire data may not scale well across longer time frames (and perhaps space), if our goal is to train more accurate data-driven models. Yet substantial evidence from the broader computer vision literature indicates that developing larger training datasets is crucial to improve modeling. Therefore, an important emerging challenge for the wildfire modeling community may be to develop means of harmonizing, and/or otherwise effectively utilizing diverse historical wildfire data to train data-driven models. We hypothesize that developing such methods may be key to obtaining better accuracy on the WSTS+ benchmark, and therefore that the benchmark will be a useful resource for developing solutions.
@@
Figure 6: Performance breakdown by test year. Blue bars represent models trained on the original WSTS data, while red bars represent those trained on WSTS+. The bolded x-axis ticks highlight original test years from WSTS.
@@
Figure 7: UMAP visualization of the inputs features across years. Each point represents the encoded features at the deepest layer of our best Res18-Unet encoder, with blue indicating original WSTS year and red newly added years in WSTS+

### v3

@@ 5.2 Time-Series Input, $T=5$
Existing models for the time-series scenario generally adopt one of two approaches: (i) a data-level fusion, or (ii) a feature-level fusion. In data-level fusion, the features for each day, $\tilde{x}(t)\in\mathbb{R}^{H\times W\times C}$ are concatenated along the feature dimension into one input tensor, $x(t)=[\tilde{x}(t-1)|,...,|\tilde{x}(t-T)]\in\mathbb{R}^{H\times W\times CT}$, after which they can be processed in the same manner as single-day input (see Sec. 2 for problem notation). Therefore, we adopt our best-performing $T=1$ models from Sec. 5.1, and their hyperparameter settings, and evaluate them for data-level fusion. As a reference, we also include the reported results of the Res18-Unet [gerard2023wildfirespreadts] when it was applied for data-level fusion.
In this context, feature-level fusion implies that we use a shared encoder to first extract features (or embeddings) independently for each day of our input, $\tilde{z}(t)=f_{\theta_{En}}(\tilde{x}(t))$ so that we have a collection of features, $z(t)=\{\tilde{z}(t-i)\}_{i=1}^{T}$, which are utilized as input into a subsequent model for joint processing (i.e., fusion). The current SOTA accuracy on WSTS, both for the time-series setting, and overall, was obtained with a UTAE model [garnot2021panoptic], as reported in [gerard2023wildfirespreadts]. Furthermore, the UTAE achieved superior accuracy despite having just 1.1M parameters - significantly fewer than many other models considered (e.g., the Res18-Unet has 14.3M). Therefore, we focus our modeling improvements on the UTAE from [gerard2023wildfirespreadts].
@@
We develop two improved UTAE models, referred to as UTAE and UTAE(Res18). Sec. 8.1 introduces key notation and background supporting the design of each model below, and the new positional encodings.
Our UTAE Model. Our UTAE includes two major improvements over the UTAE[gerard2023wildfirespreadts]. The first improvement is to adopt all of the changes investigated for the single-day models from Sec. 5.1. Pursuant to this, following previous work convention, we did a joint search over the following hyperparameters using a single fold of the WSTS benchmark: pre-training (or not), learning rates ([1e2,1e3,1e4,1e5][1e-2,1e-3,1e-4,1e-5]), and the type of loss (Focal, BCE, Jaccard, and Dice loss). The second improvement is the introduction of a novel positional encoding in the temporal fusion utilized by the UTAE. To our knowledge, this modification is novel within the vision and wildfire literature. Specifically, instead of using day-of-year positional encodings, as done in [garnot2021panoptic, gerard2023wildfirespreadts], where $\bar{t}\in[1,365]$, we propose to use a absolute positional encoding that indicates the relative position of each day’s set of features within the time-series, so that $\bar{t}\in[1,...,T]$ for a T-day input. We hypothesize that the features (especially the fire mask) from the most recent day of the fire will be most important for making predictions, and therefore, this relative position information will be much more important than its position in the year. Furthermore, it may be difficult for the models to infer relative positional information from day-of-year encodings, potentially undermining performance.
Our UTAE(Res18) Model. This model is obtained by making one additional improvement to our UTAE model. The encoder utilized in the UTAE[gerard2023wildfirespreadts] is relatively small (in terms of free parameters). Therefore, in a similar fashion to our investigation in Sec. 5.1, we replace the existing UTAE’s encoder with a pre-trained ResNet-18.
@@
Tab. 2 reports the accuracy (in terms of AP) of our time-series models on the WSTS benchmark, categorized by the type of fusion performed: data-level or feature-level. Regarding data-level fusion, our Res18-Unet, Swin-Unet, and SegFormer all substantially outperform the existing Res18-Unet[gerard2023wildfirespreadts] across all combinations of input features, with the Res18-Unet providing the best overall AP (AP=0.472, on Vegetation features). Regarding feature-level fusion, our two UTAE models (UTAE and UTAE(Res18)) substantially outperform the existing UTAE[gerard2023wildfirespreadts], which is the current SOTA model on WSTS, both for time-series input ($T>1$) and overall. Our UTAE(Res18) model achieves the highest overall performance for each combination of input features, across both single-day and time-series models. In particular, our UTAE(Res18) achieves the highest overall AP with the Vegetation (Veg) feature subset, leading to a new overall SOTA performance on WSTS of AP=0.478.
Notably, our results indicate that models receiving time-series input generally outperform those with single-day input. This is especially apparent when comparing data-level fusion models, such as Res18-Unet and SwinUnet, with their single-day counterparts, since they have few architectural differences. Most existing wildfire spread prediction in the literature has focused on the single-day input; however, our findings here corroborate those from [gerard2023wildfirespreadts] and suggest that time-series modeling is a promising emerging modeling strategy.
Our results also provide evidence that each modeling change is beneficial. As discussed, our UTAE included several applicable improvements discussed for our single-day models in Sec. 5.1, as well as our improved temporal encodings described in this sub-section. We therefore conducted an ablation experiment, reported in Tab. 3, to demonstrate that our modified positional encodings provide additional benefits. To show that the pre-trained ResNet-18 encoder is beneficial, we can compare the performance of UTAE(Res18) and UTAE in Tab. 2: the pre-trained ResNet-18 is the only difference between these two models.
Table 3: Test AP of UTAE trained on Vegetation features using the original Absolute positional encodings from [gerard2023wildfirespreadts], versus our proposed Relative positional encodings
@@
Finally, we observe that increasing the number of input features is not always beneficial, which is consistent with [gerard2023wildfirespreadts], where the best AP was often achieved with the Veg or Multi feature sets rather than the All set. This suggests that the explanatory power of some features is outweighed by the cost of (often significantly) increasing the input dimensionality. We hypothesize that this may be due to the low resolution and/or noise present in some features, such as the weather forecast features, which have a resolution of 27 km, while the fire masks have a resolution of 375 m.
@@ 6 The WSTS+ Benchmark
Our results on the WSTS benchmark indicated that relatively simple models performed best, such as those based upon a ResNet-18, rather than models utilizing larger encoders (e.g., ResNet-50) or those utilizing attention (e.g., SwinUnet). This contrasts sharply with the broader vision literature where larger models tend to perform best, given sufficient quantities of training data. Therefore, we hypothesize that collecting more training data would facilitate the use of larger models, yielding superior modeling performance. To investigate this hypothesis, we expand the original WSTS benchmark by curating four additional years of historical wildfire data: 2016, 2017, 2022, and 2023. Our extended dataset, termed WSTS+, contains twice the number of years of historical wildfire data, expands the geographic diversity of the benchmark, and is – to our knowledge – the largest public benchmark for time-series next-day wildifre spread prediction. We visualize the geographic distribution of WSTS+ events in Fig. 3 and find that it much of the new data is in the Western United States, similar to WSTS, but that it includes some unique locations there, and some additional data in the eastern states. Tab. 4 summarizes the differences between both datasets in terms of numbers of years, fire events, total images, and active fire pixels. Further collection details can be found in Sec. 10.1 of the supplement.
@@
Years 4 (2018-2021) 8 (2016-2023) +100+100
Fire Events 607 1,005 +65.6+65.6
Total Images 13,607 24,462 +79.8+79.8
Active Fire Px 1,878,679 2,638,537 +40.4+40.4
@@ 6.1 Benchmarking Models with WSTS+
As compared to WSTS, we propose a new scheme for evaluating models using WSTS+, which exploits its greater size to significantly reduce computational complexity compared to WSTS’s 12-fold cross-validation – thereby making the benchmark more accessible to researchers – while maintaining a similar level of real-world rigor. For WSTS+, we propose to divide the available data into four folds that each contain two consecutive years of historical wildfire data. We then evaluate models using four-fold cross-validation, where in each iteration, one fold of data is used for testing, one fold for validation, and two folds for training, as illustrated in Fig. 4. To ensure that the testing and validation sets have the same relative temporal distance to the training set, we always select them so that they are non-consecutive. This results in four-fold cross-validation instead of the twelve-fold cross-validation utilized in WSTS, making it far less computationally intensive. At the same time, this approach doubles the quantity of data in the training and validation sets, ideally allowing researchers to train larger and more sophisticated models. Lastly, because two consecutive years of data are included in the test set, the benchmark still evaluates models under challenging realistic testing conditions.
@@
UTAE(Res18) 5 $\mathbf{0.354\pm 0.113}$ $\mathbf{0.363\pm 0.129}$ $0.350\pm 0.117$}
@@ 6.2 Experimental Results with WSTS+
Using our updated cross-validation scheme, we train our best $T=1$ models and our best $T=5$ model on WSTS+ and report the results in terms of mean average precision across all three feature sets in Tab. 5. We see that the performance rank-order of our three models is still similar on WSTS+ as compared to WSTS. However, the overall performance is significantly lower for these models on WSTS+ as compared to WSTS (by roughly 0.1 AP). These results seem to contradict our initial hypothesis that additional training data would enable larger models and improve accuracy. To investigate further, Fig. 5 reports the per-year performance for a Res18-Unet trained on either WSTS or WSTS+ (denoted Res18-Unet(WSTS) and Res18-Unet(WSTS+), respectively; see caption for details). These results reveal that both models obtain very similar AP on every testing year, despite Res18-Unet(WSTS+) being trained on twice as many years of data in each fold as Res18-Unet(WSTS). Since the two models perform similarly across all years, the lower overall performance obtained on the WSTS+ benchmark in Tab. 5 is likely due to the greater apparent difficulty of the new testing years (2016, 2017, 2022, and 2023), rather than lower predictive accuracy of the models trained on WSTS+.
@@
Figure 5: Performance breakdown by test year. Blue bars represent models trained on the original WSTS data, termed Res18-Unet(WSTS), while red bars represent those trained on WSTS+, termed Res18-Unet(WSTS+). The bolded x-axis ticks highlight original test years from WSTS. For Res18-Unet(WSTS+), we stratify its performance by year. For Res18-Unet(WSTS), we stratify by year to obtain performance for 2018 to 2021. To obtain performance on the remaining years, we select the cross-validation fold with the best-performing model (as judged by its test fold error) and report its performance on the newly added WSTS+ years.
@@
Figure 6: UMAP visualization of the input features across years. Each point represents the encoded features at the deepest layer of our best Res18-Unet encoder, with blue indicating original WSTS year and red newly added years in WSTS+
6.3 Domain Shift: A Potential Challenge to Scaling Data-Driven Wildfire Modeling
The results in Fig. 5 raise a question: why does significantly increasing the quantity and diversity of the training data in WSTS+ lead to little or no improvement? In Sec. 10.2 of the supplement, we present evidence that, in pursuit of WSTS+, we accurately reproduced the preprocessing used for WSTS. Therefore, we argue here that this result is likely caused by cross-year domain shift [quinonero-candela_dataset_2010], wherein the joint probability distribution of the features and targets, denoted $p(x(t),y(t))$, varies across years. There is a large literature about identifying and addressing domain shift (e.g., see [quinonero-candela_dataset_2010, pan2009survey]), and comprehensively addressing these problems is beyond the scope of this work. However, we seek here to provide evidence that domain shift is present in historical wildfire datasets, provide an initial characterization of it, and discuss the implications of it. There are many different types of domain shift based upon precisely how $p(x(t),y(t))$ changes from training to testing conditions (or across years in our case) [quinonero-candela_dataset_2010, kouw_introduction_2019, kull_patterns_nodate]). We consider here two widely-studied types of shift: concept and covariate shift.
Concept Shift refers to changes in the conditional distribution $p(y(t)|x(t))$, which in a regression context generally implies that the true underlying function $y(t)=f(x(t))$ is changing. Under this hypothesis, we would expect that combining multiple years of data would likely lead to significant reductions in accuracy, since each year exhibits a different underlying relationship. Our results in Fig. 5 indicate that adding two additional years of training data, as is done in WSTS+, did not significantly impact accuracy, suggesting that significant concept shifts are unlikely. In Tab. 6 we also report the results of an experiment where we train eight Res18-Unet models - one on data from each year - and then test each model on a disjoint test set from each year (see Sec. 10.4 of the supplement for experimental details). The results indicate that, given a specific single testing year, most models achieve relatively similar accuracy, also suggesting they are each learning similar concepts.
Covariate Shift refers to change in the marginal distribution, $p(x(t))$. Covariate shifts are thought to often have limited negative impact for high-capacity models [quinonero-candela_dataset_2010], such as our DNNs here. Under this hypothesis, additional years of training data may be either beneficial or neutral, but not especially detrimental. This hypothesis is therefore consistent with Fig. 5. It is also corroborated by the results of Tab. 6 if we note that, for a specific testing year, the accuracy of most single-year models is similar to that obtained by WSTS models (trained on two years) and WSTS+ models (trained on four years) in Fig. 5. For example, if we take 2018 as the testing year, then the average of single-year models in Tab. 6 is nearly the same as the WSTS and WSTS+ models in Fig. 5. In other words, despite significant differences in the years included in, and total size of the training data, these models usually perform similarly. Notably, this also indicates that the WSTS benchmark did not benefit from additional training data: the WSTS models in Fig. 5 do not perform differently (on average) than the single-year models in Tab. 6. As additional quantitative evidence, Sec. 10.3 in the supplement provides substantial additional evidence that there is significant inter-year variability in environmental conditions (e.g., landcover composition, vegetation indices, and weather variables) and fire size and behavior. This is summarized and corroborated by Fig. 6, which presents a UMAP visualization of the features extracted by our Res18-Unet for each year in WSTS+. The results show that there is significant overlap in the feature distributions, but there are also significant apparent shifts across years.
Conclusions It is well-known within the fire science community that fire spread is impacted by a diverse set of environmental factors (e.g., weather, topography, and fuel) [rothermel1972mathematical, holsinger2016weather], and these factors vary substantially across space and time. For example, fire behavior experts have long established that annual weather patterns strongly influence both fire prevalence and extent [swetnam1990fire]. Most of these important environmental factors are encoded by one or multiple input features in the WSTS dataset, providing a plausible physical basis for inter-year covariate shift. Our analysis above provides substantial additional evidence that there is significant inter-year covariate shift in historical wildfire data, which would explain the limited benefits of additional training data in WSTS+.
Table 6: Cross-year results: Rows show the year the model was trained on, while columns show the year the model was tested on.
Year 2016 2017 2018 2019 2020 2021 2022 2023 Avg
2016 0.350 0.291 0.490 0.276 0.173 0.544 0.268 0.416 0.351
2017 0.242 0.300 0.487 0.288 0.180 0.568 0.301 0.437 0.351
2018 0.265 0.297 0.576 0.313 0.194 0.595 0.344 0.465 0.381
2019 0.219 0.259 0.455 0.329 0.159 0.530 0.324 0.428 0.338
2020 0.222 0.263 0.501 0.285 0.220 0.572 0.295 0.460 0.352
2021 0.253 0.321 0.534 0.330 0.187 0.649 0.328 0.465 0.384
2022 0.227 0.249 0.460 0.261 0.163 0.508 0.390 0.416 0.334
2023 0.242 0.279 0.483 0.289 0.157 0.568 0.324 0.582 0.365
Avg 0.253 0.282 0.498 0.296 0.179 0.567 0.322 0.459 0.357